# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ IV.4.3 Machine Learning Based Kinetic Energy Density Functional
In conclusion, the MPN KEDF exhibits promising potential in accurately predicting the energies of complex alloy systems. The high accuracy achieved by the MPN KEDF suggests its capability to provide reliable energy predictions for a wide range of materials.
V Methods in Numerical Atomic Orbital Basis
V.1 Numerical Atomic Orbitals
The efficiency and accuracy first-principles are largely determined by the basis sets. Although the accuracy of plane-wave basis mentioned in Sec. IV can be systematically improved by increasing the kinetic energy cutoff, its computational costs are formidable for systems of hundreds or thousands of atoms. On the other hand, the atomic orbital basis sets often provide more efficient description and room for optimization (like linear-scaling methods) at the cost of minor accuracy loss.
In the past two decades, numerical atomic orbitals (NAOs) have become a competitive option for basis set types, especially for simulating large systems. The NAO basis sets are favorable for their flexibility, excellent balance between accuracy and efficiency, the capability of supporting linear scaling algorithms via their strict locality, etc. Over the past several decades, several methods have been proposedTalman (1978); Sharafeddin et al. (1992); Toyoda and Ozaki (2010); Cerioni et al. (2012); Soler et al. (2002); Blum et al. (2009) to perform numerical integrations for NAOs more efficiently, thus efficient computation of large systems on the NAO basis can be achieved.
Unlike plane-wave basis, there is no unique way to construct NAOs. For example, a popular class of methods involves solving isolated atoms subject to confining potentialsSankey and Niklewski (1989); Porezag et al. (1995); Horsfield (1997); Junquera et al. (2001); Soler et al. (2002). Based on confining potentials, Blum et alBlum et al. (2009) proposed to generate NAOs by iteratively picking up basis functions one by one from a pool of predefined candidates, to seek the best improvement of a target energy. Alternatively, OzakiOzaki (2003); Ozaki and Kino (2004) suggests that eigenfunctions of isolated atoms in confining potentials serve as “primitive orbitals” in terms of which NAOs are expanded, and expansion coefficients can be optimized alongside self-consistent cycles.
Apart from the above energy-based methods, another class of methods constructs NAOs towards some reference states. For example, Sanchez-Portal et alSanchez-Portal, Artacho, and Soler (1995); Sánchez-Portal, Artacho, and Soler (1996)proposed to optimize the following “spillage”
@@ V.1 Numerical Atomic Orbitals
$=\sum_{n\mathbf{k}}\left\|(1-\hat{P}(\mathbf{k}))\ket{\psi_{n}(\mathbf{k})}\right\|^{2},$ (88)
where $\{\psi_{n}(\mathbf{k})\}$ denotes reference states from plane-wave calculations, $\hat{P}(\mathbf{k})$ is the projection operator onto the subspace of Bloch functions characterized by $\mathbf{k}$ and spanned by atomic orbitals:
$P({\mathbf{k}})\equiv\sum_{\mu\nu}\ket{\phi_{\mu}(\mathbf{k})}S_{\mu\nu}^{-1}(\mathbf{k})\bra{\phi_{\nu}(\mathbf{k})}$ | (89)
and $S_{\mu\nu}(\mathbf{k})\equiv\innerproduct{\phi_{\mu}(\mathbf{k})}{\phi_{\nu}(\mathbf{k})}$ is the overlap matrix. In their original scheme, NAOs are chosen to be combinations of pseudo-atomic orbitals (eigenfunctions of isolated atoms with pseudo-potentials) or Slater-type orbitals, and reference systems are ordinary solids. Based on the spillage formalism, Chen, Guo, and He (CGH)Chen, Guo, and He (2010a) proposed to construct NAOs with localized spherical wavesHaynes and Payne (1997), and reference systems are chosen to be a series of isolated dimers/trimers of variable bond lengths. The CGH basis set turns out to have a very small eggbox effectLi et al. (2016) and has been successfully applied to several studiesLiu et al. (2017); Liu, Ren, and He (2019); Wang et al. (2019b). Recently, Lin, Ren, and He (LRH)Lin, Ren, and He (2021b) show that the CGH basis set can be further improved by introducing a gradient term to the original spillage ($\hat{p}$ is the momentum operator):
$\mathcal{S}^{\text{LRH}}=\mathcal{S}+\sum_{n\mathbf{k}}\left\|\hat{p}(1-\hat{P}(\mathbf{k}))\ket{\psi_{n}(\mathbf{k})}\right\|^{2}.$ (90)
CGH and LRH basis sets paired with SG15 pseudopotentialsSchlipf and Gygi (2015) are publicly availableABACUS team (2020) and have been extensively tested (see Sec. V.2). In ABACUS, the implementation of NAO basis relies on CGH and LRH basis.
V.2 Kohn-Sham Equation in NAO Basis
@@
$H(\mathbf{k})C(\mathbf{k})=S(\mathbf{k})C(\mathbf{k})E(\mathbf{k}),$ (91)
in which $H(\mathbf{k})$, $C(\mathbf{k})$, and $S(\mathbf{k})$ are the Hamiltonian matrix, the electronic wave function coefficients of NAOs, and overlap matrix, respectively. The $E(\mathbf{k})$ is a diagonal matrix with KS eigenvalues.
@@
in which functions $\phi$ (basis or any "projector") centered at atoms spaced by $\mathbf{R}$ are distinguished by $\mu$ and $\nu$. The overlap matrix $S_{\mu\nu}(\mathbf{R})$ and the kinetic energy matrix $T_{\mu\nu}(\mathbf{R})$ are evaluated directly with this form, while the representation of the non-local part of pseudopotential requires the calculation of
@@
Equation 91 is solved independently for each $\mathbf{k}$. The representation of Hamiltonian or any operator within the momentum space is obtained via a "folding" operation
@@
in which $\mathbf{R}$ always runs over all valid neighboring cells. Band energy can be obtained through the expectation of the Hamiltonian:
$E_{\rm band}=\sum_{n\mathbf{k}}f_{n\mathbf{k}}\langle\psi_{n\mathbf{k}}|\hat{H}|\psi_{n\mathbf{k}}\rangle=\sum_{\mu\nu,\mathbf{k}}\rho_{\nu\mu}(\mathbf{k})H_{\mu\nu}(\mathbf{k})=\rm Tr(\rho H),$ (95)
in which the density matrix is defined as
$\rho_{\mu\nu}(\mathbf{k})=\sum_{n}f_{n\mathbf{k}}C_{n\mu}(\mathbf{k})C^{*}_{n\nu}(\mathbf{k}).$ (96)
The spinor index is omitted here. Real-space density matrix $\rho_{\mu\nu}(\mathbf{R})$ is calculated through Fourier transform accordingly.
V.3 Forces and Stresses
To obtain the analytical expression for force and stress, one must start with the analytical expression for energy
$E_{\rm total}$ $=\sum_{\mu\nu,\mathbf{R}}\rho_{\nu\mu}(\mathbf{R})\Big{[}T_{\mu\nu}(\mathbf{R})+V^{NL}_{\mu\nu}(\mathbf{R})\Big{]}+\int V_{\rm local}(\mathbf{r})\rho(\mathbf{r})\mathrm{d}\mathbf{r}$ (97)
$+\frac{1}{2}\iint\frac{\rho(\mathbf{r})\rho\left(\mathbf{r}^{\prime}\right)}{\left|\mathbf{r}-\mathbf{r}^{\prime}\right|}\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{r}^{\prime}+\int\epsilon_{\rm xc}(\mathbf{r})\rho(\mathbf{r})\mathrm{d}\mathbf{r}+E_{\rm II},$
where the ionic energy is given by the Ewald method Martin (2020). It should be noted that ABACUS uses Eq. 36 to calculate the total energy in practical computations. For more information about evaluating force and stress terms in ABACUS, we refer the readers to a recent review work in Ref. Lin et al., 2024.
The atomic forces are obtained by direct differentiation with respect to atomic positions
$\bm{F}_{I}$ $=-\sum_{\mu\nu}\sum_{\mathbf{R}}\rho_{\nu\mu}(\mathbf{R})\frac{\partial H_{\mu\nu}(\mathbf{R})}{\partial\mathbf{R}_{I}}-\sum_{\mu\nu}\frac{\partial\rho_{\nu\mu}(\mathbf{R})}{\partial\mathbf{R}_{I}}H_{\mu\nu}(\mathbf{R})$
$=\bm{F}_{I}^{\rm HF}+\bm{F}_{I}^{\rm Pulay}+\bm{F}_{I}^{\rm Orth}+\bm{F}_{I}^{\rm Ewald}.$ (98)
According to the Hellmann-Feynman theorem Hellmann (1935); Feynman (1939), the HF force $\bm{F}_{I}^{\rm HF}$ describes the contribution of the partial derivatives of the operator. However, since the positions of the NAOs basis set change with the atomic positions, the partial derivatives of the basis set with respect to nuclear positions also contribute to the atomic forces, which are referred to as the Pulay forces $\bm{F}_{I}^{\rm Pulay}$. The orthogonal force $\bm{F}_{I}^{\rm Orth}$ is a correction term. The final term $\bm{F}_{I}^{\rm Ewald}$ represents the contribution of ionic interactions to atomic forces.
The stress tensor can be defined as the positive derivative of the total energy Eq.97 with respect to the strain tensor (Eq. 25). Taking advantage of the relation $\frac{\partial r_{\gamma}}{\partial\epsilon_{\alpha\beta}}=\delta_{\gamma\alpha}r_{\beta}$, the stress calculation can be derived in a form very similar to that of force calculation through Lin et al. (2024)
$\frac{\partial H_{\mu\nu}}{\partial\epsilon_{\alpha\beta}}=\sum_{\gamma}\frac{\partial H_{\mu\nu}}{\partial\mathbf{R}^{\gamma}_{\mu\nu}}\frac{\partial\mathbf{R}^{\gamma}_{\mu\nu}}{\partial\epsilon_{\alpha\beta}}=\frac{\partial H_{\mu\nu}}{\partial\mathbf{R}^{\alpha}_{\mu\nu}}\mathbf{R}^{\beta}_{\mu\nu},$ (99)
where $\gamma$ represent coordinate components x, y, z. In fact, atomic force and stress are computed at almost the same place in the code. We will not elaborate further here.
We verify the accuracy of the NAOs by comparing the results with the plane-wave method. We have generated a series of NAOs, including Single-$\zeta$ (SZ), Double-$\zeta$ plus polarization functions (DZP), and Triple-$\zeta$ plus double polarization functions (TZDP) with different cutoffs based on the Pseudo-Dojo v0.4 (3plus) pseudopotential for Yb and carried out self-consistent field calculations using these orbitals. The results are in Fig. 18. It is evident that including more atomic orbitals and utilizing larger cutoffs can effectively reduce the difference in total energy between NAO and PW calculations. For example, the total energy calculated using the TZDP orbitals with a cutoff of 10 a.u. is closest to the PW results, with a deviation of about 0.0053 eV/atom.
@@ V.3 Forces and Stresses
Figure 18: (a) Comparison on energies calculated with NAO and PW basis set for the $\rm Li_{12}\rm Yb_{4}\rm Cl_{24}$ system calculated using Yb SZ/DZP/TZDP orbitals with different orbital cutoffs. (b) MAE is the mean absolute error of the energies between NAO and PW for different volumes, with the unit being eV/atom.
We have performed the finite difference tests for atomic forces using body-centered cubic (BCC) Fe (perturbing the position of one Fe atom so that the structure can deviate from the equilibrium position), which specifically involved moving one Fe atom in the positive and negative direction of the $x$-axis by a particular step size and calculating the energy. Fig. 19(a) shows the atomic forces calculated by the finite difference method (step size is 0.02 Bohr) and the analytical values at different positions of the Fe atom. The difference between them is plotted in the zoom-in figure. It is apparent that for the Fe systems, this force discrepancy is around 0.002 eV/Å. This discrepancy can be affected by a few factors, including the precision of the finite difference method (step size) and the precision of the energy (including the accuracy of the SCF calculations and numerical errors). Fig. 19(b) shows the value of this discrepancy for different step sizes used in finite difference. As the differential step size decreases, the discrepancy gradually reduces and tends toward zero, indicating that the force in ABACUS LCAO implementation is sufficiently consistent with the energy. Stress can be verified similarly by slightly changing the lattice vectors.
@@
Figure 19: By using the NAO basis set, the finite difference (FD) tests of atomic forces on the $x$-axis of BCC Fe. a) The analytical forces and the forces calculated by the finite difference method along the x-direction for different configurations. The zoom-in figure shows the difference between them; b) The deviations between FD values and analytical values along the Fe $x$-direction for different step sizes used in the finite difference method.
V.4 Diagonalization-Based Eigensolvers
In the SCF calculations using the NAO basis set of ABACUS, the combined computational time for grid integrals and generalized eigenvalue solving exceeds 90%. For the generalized eigenvalue and eigenvector solving, ABACUS invokes the libraries scalapack sca (2024) and ELPA elp (2024).
We have accelerated these computational loads using GPUs. ABACUS integrates GPU-supported libraries such as ELPA Yu et al. (2021), cuSolvermpTal et al. (2024), and cuSolver. Although cuSolvermp demonstrates superior diagonalization performance on NVIDIA GPU clustersTal et al. (2024), we still recommend users to utilize ELPA as the diagonalization library for large-scale, multi-node, multi-GPU computations in most cases. This recommendation is based on three key factors related to performance and ease of use: 1) The generalized eigenvalue solving process in ABACUS using the NAO basis set does not require solving for all eigenvectors, and ELPA supports configuring the number of eigenvectors to be solved; 2) The self-consistent iterative calculations in ABACUS involve solving eigenvalues and eigenvectors multiple times, and ELPA can avoid repeated S-matrix decompositions; 3) ELPA provides broad support for various GPU computing stacks, including those from NVIDIA, AMD, and INTEL, reducing the barrier for users across different hardware platforms. Besides, ABACUS also provides a low-scaling solver PEXSI (see Sec. VII.8).
V.5 Hybrid Funtional
Hybrid density functionals (HDFs) formulated within the generalized KS frameworkSeidl et al. (1996) belongs to the fourth rung of the Jacob’s ladderPerdew and Schmidt (2001). It overcomes the drawbacks of the (semi-)local XC functionals, namely the self-interaction errorsPerdew and Zunger (1981); Mori-Sánchez, Cohen, and Yang (2008) related to the underestimation of the band gaps.
ABACUS supports pure exact exchange (EXX), namely Hartree-Fock (HF). Based on that, we implement various hybrid functionals, including HSEHeyd, Scuseria, and Ernzerhof (2003), PBE0Ernzerhof and Scuseria (1999); Adamo and Barone (1999) and SCAN0Hui and Chai (2016), which are mixtures of EXX and (semi-)local XC functionals. The local part is calculated with LibXC Lehtola et al. (2018), while the exact (or Fock) exchange energy is given by the 2-electron integrals of the Kohn-Sham orbitals:

### v3

@@ IV.1 Numerical Atomic Orbitals
where $\{\psi_{n}(\mathbf{k})\}$ denotes reference states from plane-wave calculations, and $\hat{P}(\mathbf{k})$ is a projection operator defined as
$|.\displaystyle\hat{P}({\mathbf{k}})=\sum_{\mu\nu}\ket{\phi_{\mu}(\mathbf{k})}S_{\mu\nu}^{-1}(\mathbf{k})\bra{\phi_{\nu}(\mathbf{k})}$. (80)
Here $S_{\mu\nu}(\mathbf{k})$ is the overlap matrix. In the original work, NAOs are chosen to be combinations of pseudo-atomic orbitals (eigenfunctions of isolated atoms with pseudo-potentials) or Slater-type orbitals. Based on the spillage formalism, Chen et al. Chen, Guo, and He (2010a) proposed to construct NAOs with localized spherical Bessel functions as basis Haynes and Payne (1997), and reference systems are chosen to be a series of isolated dimers/trimers of variable bond lengths. Recently, Lin et al. Lin, Ren, and He (2021b) further improved the basis by introducing a gradient term to the original spillage
@@
where $\hat{p}$ is the momentum operator. To use the NAO basis set in ABACUS, the collection of NAO basis sets paired with SG15 pseudopotentialsSchlipf and Gygi (2015) is publicly availableABACUS team (2020).
IV.2 Kohn-Sham Equation in NAO Basis
@@
$\bm{H}(\mathbf{k})\bm{C}(\mathbf{k})=\bm{S}(\mathbf{k})\bm{C}(\mathbf{k})\bm{E}(\mathbf{k}),$ (82)
in which $\bm{H}(\mathbf{k})$, $\bm{C}(\mathbf{k})$, and $\bm{S}(\mathbf{k})$ are the Hamiltonian matrix, the electronic wave function coefficients of NAOs, and the overlap matrix, respectively. The $\bm{E}(\mathbf{k})$ is a diagonal matrix with KS eigenvalues. The overlap matrix element $S_{\mu\nu}(\mathbf{k})$ is given by
$S_{\mu\nu}(\mathbf{k})=\sum_{\mathbf{R}}\langle\phi_{\mu 0}|\phi_{\nu\mathbf{R}}\rangle e^{i\mathbf{k}\cdot\mathbf{R}},$ (83)
where $\phi_{\nu\mathbf{R}}(\mathbf{r})$ is a NAO centered in the cell with lattice vector $\mathbf{R}$.
@@
where functions $\phi$ (basis or any "projector") centered at atoms spaced by $\mathbf{R}$ are distinguished by $\mu$ and $\nu$. The overlap matrix $S_{\mu\nu}(\mathbf{R})$ and the kinetic energy matrix $T_{\mu\nu}(\mathbf{R})$ are evaluated directly with this form, while the representation of the non-local part of pseudopotential requires the calculation of
@@
Here, indices $i$ and $j$ run over all projects for atom $I$. Equation 82 is solved independently for each $\mathbf{k}$ point. The representation of Hamiltonian or any operator within the momentum space is obtained via a "folding" operation
@@
in which $\mathbf{R}$ always runs over all valid neighboring cells. After diagonalization of the Hamiltonian matrix, the eigenvalues and corresponding Kohn-Sham wave functions can be obtained.
To solve the generalized eigenvalue problem in Eq. 82, ABACUS supports numerical libraries such as ScaLAPACK sca (2024) and the Eigenvalue SoLvers for Petaflop Applications (ELPA) elp (2024). ABACUS also provides the pole expansion and selected inversion (PEXSI) method as a low-scaling alternative (see Sec. VI.8 for details). Additionally, we have implemented GPU acceleration through specialized numerical libraries, such as the GPU-enabled ELPA implementation Yu et al. (2021). After diagonalization, the density matrix can be constructed via the formula
$\rho_{\mu\nu}(\mathbf{R})=\frac{1}{N_{\mathbf{k}}}\sum_{n\mathbf{k}}f_{n\mathbf{k}}c_{n\mu,\mathbf{k}}c_{n\nu,\mathbf{k}}^{\ast}e^{-i\mathbf{k}\cdot\mathbf{R}},$ (87)
where $f_{n\mathbf{k}}$ is the occupation, and the electronic wave functions are $\{c_{n\mu,\mathbf{k}}\}$. Subsequently, the electron charge density $\rho(\mathbf{r})$ can be evaluated through real-space grids.
@@
Figure 17: (a) Comparison on energies calculated with NAO and PW basis sets for the $\rm Li_{12}\rm Yb_{4}\rm Cl_{24}$ system calculated using Yb SZ/DZP/TZDP orbitals with different orbital cutoffs. (b) Mean absolute error (MAE, in eV/atom) of the energies between NAO and PW basis sets for different volumes.
To assess the accuracy of the NAO basis set, we conducted a systematic comparison with PW results. A series of NAO basis sets—including Single-$\zeta$ (SZ), Double-$\zeta$ plus polarization functions (DZP), and Triple-$\zeta$ plus double polarization functions (TZDP)—were generated with varying cutoff radii based on the Pseudo-Dojo v0.4 (3plus) pseudopotential. We then performed SCF calculations using these basis sets, and the resulting total energies are presented in Fig. 17. Our analysis demonstrates that incorporating a larger number of atomic orbitals and employing higher cutoff values significantly reduces the discrepancy between NAO and PW results. Notably, the TZDP basis set with a cutoff of 10 a.u. yields the closest agreement with the PW results, exhibiting a minimal deviation of approximately 0.0053 eV/atom.
IV.3 Forces and Stress
To obtain the analytical expression for atomic forces and stress, one must start with the analytical expression for the total energy
$E_{\rm tot}$ $=\sum_{\mu\nu\mathbf{R}}\rho_{\nu\mu}(\mathbf{R})\Big{[}T_{\mu\nu}(\mathbf{R})+V^{\rm NL}_{\mu\nu}(\mathbf{R})\Big{]}+\int V_{\rm local}(\mathbf{r})\rho(\mathbf{r})\mathrm{d}\mathbf{r}$ (88)
$+\frac{1}{2}\iint\frac{\rho(\mathbf{r})\rho\left(\mathbf{r}^{\prime}\right)}{\left|\mathbf{r}-\mathbf{r}^{\prime}\right|}\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{r}^{\prime}+E_{\mathrm{xc}}[\rho(\mathbf{r})]+E_{\rm II},$
where the kinetic and nonlocal pseudopotential matrices are respectively labeled as $T_{\mu\nu}(\mathbf{R})$ and $V_{\mu\nu}^{\rm NL}(\mathbf{R})$, while the local pseudopotential is $V_{\rm local}$ and the ionic energy $E_{\rm II}$ is given by the Ewald method Martin (2020). It should be noted that ABACUS uses Eq. 33 to calculate the total energy in practical computations. For more information about evaluating force and stress terms in ABACUS, we refer the readers to a recent review work in Ref. Lin et al., 2024.
The atomic forces are obtained by direct differentiation of the total energy $E_{\rm tot}$ with respect to atomic positions
$\bm{F}_{I}$ $=-\sum_{\mu\nu\mathbf{R}}\rho_{\nu\mu}(\mathbf{R})\frac{\partial H_{\mu\nu}(\mathbf{R})}{\partial\mathbf{R}_{I}}-\sum_{\mu\nu}\frac{\partial\rho_{\nu\mu}(\mathbf{R})}{\partial\mathbf{R}_{I}}H_{\mu\nu}(\mathbf{R})$
$=\bm{F}_{I}^{\rm FH}+\bm{F}_{I}^{\rm Pulay}+\bm{F}_{I}^{\rm Orth}+\bm{F}_{I}^{\rm Ewald}.$ (89)
Here we utilize the Feynman-Hellmann theorem Hellmann (1935); Feynman (1939) and divide the total forces into four components. First, the Feynman-Hellmann force $\bm{F}_{I}^{\rm FH}$ describes the contribution of the partial derivatives of the Hamiltonian operator. In this term, the kinetic and nonlocal pseudopotential terms are evaluated through the two-center integration technique, while the local pseudopotential contribution is evaluated through the plane wave basis as shown in Eq. 35. Second, the Pulay force $\bm{F}_{I}^{\rm Pulay}$ describes the partial derivatives of the NAO basis set with respect to ionic positions, as unlike the plane wave basis set, the NAO basis set changes with atomic positions. Third, the orthogonal force $\bm{F}_{I}^{\rm Orth}$ arises from the nonorthogonality of the NAO basis set, and an energy density matrix must be evaluated to yield this term. Finally, the contribution of ionic interactions to atomic forces is included in $\bm{F}_{I}^{\rm Ewald}$. The stress tensor defined in Eq. 23 can also be derived using the NAO basis set, the formulas are similar to the above force formulas. For more details, please refer to Ref. Lin et al., 2024.
To evaluate the consistency of atomic forces using the NAO basis, we performed finite-difference tests on a body-centered cubic (BCC) iron (Fe) system. Specifically, a single Fe atom was displaced along the $x$-axis in both positive and negative directions by a defined step size, and the corresponding energy changes were computed. Fig. 18(a) compares the atomic forces derived from the finite-difference method (using a step size of 0.02 Bohr) with analytically calculated forces at varying atomic positions. The inset illustrates the residual differences between the two methods. For the Fe system, the observed force discrepancy is on the order of 0.002 eV/Å. This deviation is influenced by several factors, including the precision of the finite-difference step size and the numerical accuracy of the SCF calculations. Further investigation of the step-size dependence (Fig. 18(b)) reveals that the discrepancy diminishes as the step size decreases, asymptotically approaching zero. This behavior confirms the consistency between the NAO-implemented forces and the energy landscape in ABACUS.
@@ IV.3 Forces and Stress
Figure 18: By using the NAO basis set, the finite difference (FD) tests of atomic forces on the $x$-axis of BCC Fe. (a) Analytical forces and the forces calculated by the FD method along the $x$-direction for different configurations. The zoom-in figure shows the difference between them. (b) Deviations between FD values and analytical values along the Fe $x$-direction for different step sizes used in the FD method.
IV.4 Hybrid Functional
Hybrid density functionals (HDFs) formulated within the generalized KS framework Seidl et al. (1996) belongs to the fourth rung of the Jacob’s ladder Perdew and Schmidt (2001). It overcomes the drawbacks of the local and semi-local XC functionals, namely the self-interaction errors Perdew and Zunger (1981); Mori-Sánchez, Cohen, and Yang (2008) related to the underestimation of the band gaps in solids. The ABACUS package supports pure exact exchange (EXX) calculations, specifically the Hartree-Fock (HF) method. Building upon this capability, we have implemented several widely-used hybrid functionals, including the Heyd-Scuseria-Ernzerhof (HSE) Heyd, Scuseria, and Ernzerhof (2003), PBE0  Ernzerhof and Scuseria (1999); Adamo and Barone (1999b), and SCAN0 Hui and Chai (2016). The local (semilocal) components of these functionals are computed using the LibXC library Lehtola et al. (2018), while the exact (Fock) exchange energy is evaluated through two-electron integrals of the Kohn-Sham orbitals.
$H^{\text{EXX}}_{\mu\nu}(\mathbf{R})=$ $\sum_{\lambda\sigma}\sum_{\mathbf{R}_{1}}\sum_{\mathbf{R}_{2}}\rho_{\lambda\sigma}(\mathbf{R}_{2}-\mathbf{R}_{1})(\phi^{\mathbf{0}}_{\mu}\phi^{\mathbf{R}_{1}}_{\lambda}|\phi^{\mathbf{R}_{2}}_{\sigma}\phi^{\mathbf{R}}_{\nu}),$ (90)
in which the real space density matrix $\rho_{\lambda\sigma}(\mathbf{R})$ is defined in Eq. 87 and the 4-index integrals are defined as
$(\phi^{\mathbf{0}}_{\mu}\phi^{\mathbf{R}_{1}}_{\lambda}$ $|\phi^{\mathbf{R}_{2}}_{\sigma}\phi^{\mathbf{R}}_{\nu})$ (91)
$=\iint\phi_{\mu}^{\mathbf{0}}(\mathbf{r})\phi_{\lambda}^{\mathbf{R}_{1}}(\mathbf{r})v(\mathbf{r-r^{\prime}})\phi_{\sigma}^{\mathbf{R}_{2}}(\mathbf{r^{\prime}})\phi_{\nu}^{\mathbf{R}}(\mathbf{r^{\prime}})\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{r^{\prime}},$
with $v(\mathbf{r}-\mathbf{r}^{\prime})=1/|\mathbf{r}-\mathbf{r}^{\prime}|$ for the bare Coulomb in PBE0 and HF, while $v(\mathbf{r}-\mathbf{r}^{\prime})=\mathrm{erfc(\omega|\mathbf{r}-\mathbf{r}^{\prime}|)}/|\mathbf{r}-\mathbf{r}^{\prime}|$ for short-range EXX in HSE.
In fact, the above EXX effective potential exhibits an unfavorable $O(N^{4})$ scaling behavior, making it significantly more computationally demanding than local functionals in terms of both time and memory requirements. To mitigate this challenge, several algorithmic approaches have been developed, most notably the density fitting (DF) technique Whitten (1973) and its specialized implementation for localized atomic orbitals known as resolution of identity (RI) Feyereisen, Fitzgerald, and Komornicki (1993); Vahtras, Almlöf, and Feyereisen (1993); Ren et al. (2012). These methods reduce the four-center integrals to combinations of three- and two-center integrals by expanding the orbital products using auxiliary basis functions (ABFs).
For HDF calculations in large periodic systems, the local resolution of identity (LRI) approach Levchenko et al. (2015); Ihrig et al. (2015); Lin, Ren, and He (2020) enables linear-scaling EXX calculations by exploiting the inherent locality of both atomic orbitals (AOs) and ABFs. Extensive benchmarking has demonstrated that with properly chosen ABFs, the LRI approximation achieves sufficient accuracy for HDF calculations Lin, Ren, and He (2020). Residual discrepancies in cohesive properties and band gaps primarily originate from differences in core-valence interaction treatments (pseudopotentials versus all-electron methods) rather than basis set limitations Ji et al. (2022).